Calculate the dot product of D and E, where D=7i - 3j + 2k and E=4i + 5j - 3k.
1
Expert's answer
2020-10-25T18:29:42-0400

We can calculate the dot product as follows:


"\vec{D}\cdot\vec{E}=D_x\cdot E_x+D_y\cdot E_y+D_z\cdot E_z,""\vec{D}\cdot\vec{E}=7\cdot 4+(-3)\cdot 5+2\cdot (-3)=7."

Answer:

"\vec{D}\cdot\vec{E}=7".